Perturbation theorem for nonlinear systems of difference equations Jarosław Morchało, Poznań, Poland

Keywords: Difference equations, Massera type converse theorems, Liapunov function AMS Subject classification: 39A10, 39A11 Abstract: The different types of stability and properties of perturbed nonlinear systems are discussed and several theorems are proved.